# Evaluating Task-based Effectiveness of MLLMs on Charts

### v1

@@ 4.2.3 The Impact of Visual Prompts
However, GPT-4V does not exhibit significant benefits from visual prompts in Correlation and Order tasks. These tasks often challenge GPT-4V to discern complex relationships among more than three distinct elements. In such scenarios, visual prompts may lose their specificity and lead to confusion due to the introduction of multiple new visual elements, especially in tasks like Order, where the added visual information can be misleading.
Finding-8: Adopting dynamic visual prompt strategies tailored to specific task types is essential for optimizing performance and mitigating any potential negative impacts.
Figure 9: Comparing Different Prompting Methods.
4.3 The Impact of Chain-of-Charts
The Chain-of-Thought (CoT) prompt strategy has proven effective across various scenarios [52]. The key idea of CoT is to guide the model towards producing outputs that are more coherent and logical, by mimicking the step-by-step reasoning process humans employ to solve problems. Recently, Xu et al. [11] implemented the CoT strategy for ChartQA tasks, namely ChartCoT. The key idea is to pose a series of questions to progressively guide the model in comprehending the chart’s details before it formulates an answer. However, ChartCoT struggles to ensure the accuracy of GPT-4V’s responses to guiding questions, particularly with complex charts.
Chain-of-Charts Prompts. Therefore, we introduce a novel prompting strategy, termed Chain-of-Charts, which builds on the chain-of-thought approach, as shown in Figure 1-Q3. The core of Chain-of-Charts lies in orchestrating a sequence of questions and their corresponding answers (($q_{1}$, $a_{1}$), ($q_{2}$, $a_{2}$),…($q_{m}$, $a_{m}$)) to progressively guide the model towards a deeper understanding of the chart’s details, thereby enhancing its ability to formulate accurate responses.
Experimental Settings. The testing samples for this set of experiments are the same as those used for evaluating visual prompts. In addition to Chain-of-Charts, we also evaluate two useful textual prompt strategies: Role-Play [53] and Tutorial.
Shanahan et al. [53] suggested that giving large models specific roles could enhance their performance on particular tasks Inspired by this, we assigned GPT-4V the role of a visualization expert to determine whether this specialization could improve its performance. Specifically, we craft a textual prompt: “You are an expert in ChartQA tasks with specialized skills in numerical analysis...”.
Our observations also indicate that the more detailed the prompt, the more precise and accurate GPT-4V’s responses tend to be. Therefore, we created a detailed ChartQA tutorial for GPT-4V, dubbed the “Tutorial” prompt. We structure this textual prompt in a step-by-step tutoring format: “Firstly, ... Subsequently... Next, ... Finally, summarize the information gathered.
Overall Results. Figure 5(c) reports the performance of GPT-4V. In comparison to Figure 5(a), it is clear that the Chain-of-Charts strategy has significantly improved GPT-4V’s capabilities across 10 different tasks and four basic textual prompts.
Finding-9: The Chain-of-Charts prompt have comprehensively improved the use of basic textual prompts in different tasks and chart types, demonstrating its effectiveness.
Table 6(a) shows the overall accuracy of GPT-4V on 10 low-level tasks under five prompt strategies. Overall, Chain-of-Charts leads in average accuracy across all tasks with 80.49%, outperforming ChartCoT’s accuracy of 67.55% by 12.94%. Specifically, Chain-of-Charts achieves the highest accuracy on five tasks including Reasoning, Determine Range, Order, Filter, and Find Cluster, with accuracy of 73.9%, 89.2%, 72.5%, 80.8%, and 95%. This method shows significant improvements, especially in Reasoning and Order tasks, where its accuracy surpasses other methods substantially: 73.9% in Reasoning, where others do not exceed 60%, and 72.5% in Order, where the rest fall below 40%. These tasks demand precise reasoning from GPT-4V, based on accurate identification of element coordinates and values. The Chain-of-Charts prompt framework effectively provides GPT-4V with the correct value and coordinate references, significantly aiding in the accurate positioning of different elements.
Finding-10: The Chain-of-Charts prompt supplies GPT-4V with accurate chart reference information, thereby enhancing the model’s comprehension and detailed reasoning of chart structures and elements.
@@ 4.3 The Impact of Chain-of-Charts
Table 6: The Synergistic Effect of Visual and Textual Prompts (Overall Accuracy (%))
@@
Basic Textual Prompts 41.8 47.5 50.0 62.5 72.5 28.7 50.0 70.0 96.7 90.0 59.0
Role-Play 58.6 45.0 75.0 53.8 78.3 28.7 52.5 79.2 91.7 82.5 64.5
Tutorial 40.7 52.5 40.0 63.7 84.2 36.2 55.0 73.3 94.2 80.0 61.1
ChartCoT 48.2 65.0 60.0 78.8 83.3 26.2 80.0 72.5 94.2 90.0 67.6
Chain-of-Charts (ours) 73.9 50.0 70.0 58.8 89.2 72.5 80.8 99.2 95.0 95.0 80.5
@@
Basic Textual Prompts 61.1 47.5 65.0 58.8 80.8 37.5 57.5 86.7 95.8 95.0 68.9
Role-Play 65.7 57.5 55.0 71.3 89.2 38.8 67.5 86.7 98.3 90.0 73.7
Tutorial 71.4 67.5 65.0 85.0 86.7 35.0 81.7 88.3 95.8 92.5 78.0
ChartCoT 72.1 45.0 45.0 55.0 77.5 30.0 55.8 87.5 94.2 77.5 69.2
Chain-of-Charts (ours) 79.3 62.5 65.0 67.5 92.5 81.3 78.3 97.5 96.7 95.0 83.8
4.4 The Synergistic Effect of Visual and Textual Prompts
Yet, their effectiveness wanes in complex reasoning tasks such as reasoning, anomaly detection, and sorting, as shown in Figure 5(b). Conversely, the Chain-of-Charts Prompt excels in these reasoning tasks but is less effective in the Correlation task, as indicated in Figure 5(c). This observation leads us to wonder: Could combining visual prompts with Chain-of-Charts enhance GPT-4V’s performance across a spectrum of low-level ChartQA tasks?
Experimental Settings. The test samples are the same as those used for evaluating visual prompts and Chain-of-Charts.
Overall Results. Figure 5(d) shows GPT-4V’s accuracy following the integration of Chain-of-Charts and Visual Prompt, demonstrating a clear enhancement over the outcomes depicted in Figures 5(a), (b), and (c), which demonstrates the combined strategy’s effectiveness.
Table 6 reports the performance improvements in GPT-4V after integrating various textual prompts with visual prompts. Specifically, the Chain-of-Charts with visual prompt reaches an accuracy of 83.82%, outperforming the ChartCoT with visual prompt by 14.6% and exceeding the accuracy of using Chain-of-Charts alone by 3.33%. Furthermore, this combination attained the highest accuracy in six tasks.
@@ 4.4 The Synergistic Effect of Visual and Textual Prompts
Limitations about Visual and Textual Prompts. As depicted in Figure 5(d), the integration of Chain-of-Charts and visual prompts enables GPT-4V to outperform other settings. However, the improvement over using Chain-of-Charts alone is slight. We discuss the possible reasons behind this observation.
@@
First, after carefully analyzing the experimental results, we discover that GPT-4V exhibits a certain degree of hallucination in chart understanding. For example, even if the calculation process is accurate, GPT-4V may provide answers that do not match any of the multiple-choice options, leading to incorrect results. This indicates that the model’s accuracy is significantly affected by hallucination. Moreover, we also observe that GPT-4V might output numerical information unrelated to the chart even when explicitly recognizing values, further evidencing the hallucination phenomenon in chart reading.
Finding-12: The effectiveness of using prompts to improve model performance is inherently constrained by hallucination.
Second, there is a lack of research on developing systematic and standardized visual prompts specifically designed for enhancing ChartQA with MLLMs. In our experiment, we design visual prompts based on graphical overlay strategies [34]. We believe that crafting visual prompts that align more precisely with the model’s inherent mechanisms for chart comprehension could result in significant performance improvements.
Finding-13: Developing visual prompts tailored specifically to the characteristics of ChartQA task is a promising research direction.
5 Lessons Learned
Effectiveness of Textual Prompts, Visual Prompts and Their Combination. Incorporating various prompt strategies, including textual and visual prompts, significantly impacts GPT-4V’s accuracy. Textual prompts with structured candidate answers enhance reasoning capabilities, while visual prompts enhance the chart understanding through visual attention, particularly in anomaly detection and filtering tasks.
Importance of Chart Elements and Image Quality. Alterations in chart elements and the quality of chart images influence GPT-4V’s performance. Specifically, certain modifications like larger labels or the absence of data labels can improve the model’s efficiency in specific tasks by focusing its attention on visual comparisons. However, image quality degradation, especially median blurring, negatively affects the model’s ability to process numerical values accurately.
GPT-4V’s Strengths and Weaknesses in Low-level ChartQA Tasks. GPT-4V performs well in tasks requiring direct data retrieval and basic comparisons, showing high accuracy in Query and Search task categories. However, it faces challenges in more complex reasoning, anomaly detection, and correlation tasks, indicating a need for further optimization of prompting strategies and model training to overcome these limitations.
Potential for Future Application and Development. The experiments demonstrate a promising direction for enhancing MLLMs’ performance in visual data analysis through the development of specialized prompting strategies and the careful manipulation of visual elements.
6 Limitations and Future Work
Limited Chart Types. In our experiments, we set benchmarks for the performance of GPT-4V across seven widely used chart types, providing valuable insights into the model’s capabilities in chart interpretation. However, this focus inherently excludes a range of more complex chart types, such as heatmaps, radar charts, and others, which present unique analytical challenges and opportunities for data representation.
Therefore, including a more diverse chart type, especially those with complex structure and interpretation such as heat maps and radar charts, will provide a more comprehensive perspective on ChartQA for MLLM. This extension is critical for assessing the adaptability and effectiveness of MLLM in a wider range of graph interpretation tasks.
Limited Visual Prompts Design Space. Our exploration into the effectiveness of visual prompts in facilitating ChartQA tasks with GPT-4V has shown their potential to enhance model performance. Nevertheless, our investigation into the design space of visual prompts has been preliminary, lacking a comprehensive and systematic exploration of the full spectrum of visual prompt possibilities. This limitation narrows the scope of our findings and potentially overlooks more effective visual prompt strategies that could further improve the accuracy and efficiency of MLLMs in interpreting and analyzing charts.
Future research can systematically explore the design space of visual prompts tailored to various ChartQA tasks and chart types. An interesting direction is developing algorithms that can automatically generate visual prompts from given textual prompts, specific to ChartQA tasks and chart types. This would ensure that the prompts are accurately customized to improve both model interpretability and task performance.
Lacking of Considering the Data Prompts. Our approach primarily relied on chart images, neglecting the underlying data that generated these charts. This omission could hinder the model’s ability to perform more complex analysis and reasoning based on the actual data points. Future work could explore integrating the underlying data as part of the prompt, potentially through multimodal inputs, to provide a richer context for the model’s analyses.
Without Fine-tuning MLLMs. We only use the “off-the-shelf” GPT-4V to conduct evaluation, without considering other MLLMs because GPT-4V is known as one of the best models in the visual question-answering task. In addition, we don’t perform task-specific fine-tuning because we want to benchmark GPT-4V in low-level tasks and investigate the impact of textual and visual prompts, which is orthogonal to fine-tuning the MLLMs. Future work can fine-tune MLLMs using our dataset to investigate their effectiveness.
Therefore, a promising direction is to develop a framework that includes self-correction [54], debugging, or a multi-agent approach [55]—with specialized agents for data analysis, chart understanding, and textual reasoning—that could enhance the model’s accuracy and reliability in ChartQA tasks.

### v3

@@ 3.2 Experimental Results and Findings
However, GPT-4o does not exhibit significant benefits from visual prompts in Correlation and Order tasks. These tasks often challenge GPT-4o to discern complex relationships among more than three distinct elements. In such scenarios, visual prompts may lose their specificity and lead to confusion due to the introduction of multiple new visual elements, especially in tasks like Order, where the added visual information can be misleading.
Finding-8: Different tasks require tailored visual prompts for effective chart comprehension. Using the same style of visual prompts across tasks can have a negative impact on certain tasks.
◆ Exp-5: Impact of Chart Variations
Experimental Settings.
A chart can exhibit visual differences by varying its constituent elements (e.g., data labels), as illustrated in Appendix A.2 and Figure 8(a). Intuitively, these visual variations may influence GPT-4o’s performance on low-level ChartQA tasks. To investigate this hypothesis, we explore how varying chart elements affect the performance of GPT-4o. To this end, we develop 356 visual variants of 35 base charts as test samples. These variants are associated with 17,972 textual prompts, spanning 10 low-level tasks.
Overall Results.
Figure 6: The Impact of Varying Chart Elements on GPT-4o’s Performance.
Overall, Figure 6(a) indicates that most chart variants have a minor negative impact on GPT-4o’s performance. However, the absence of data labels significantly impairs its performance across seven chart types (Figure 6(a)). This is understandable, as data labels aid GPT-4o in comprehending the underlying insights conveyed by the charts. Interestingly, GPT-4o shows a performance boost of 17.5% in anomaly detection and 5.5% in filtering tasks when data labels are not present (Figure 6(b)). This implies that data labels may sometimes hinder GPT-4o’s ability to identify anomalies and filter values effectively.
Figure 6(b) reveals that certain chart variants, such as larger x/y/data labels, positively impact GPT-4o’s performance in tasks like anomaly detection, filtering, and clustering. These tasks inherently involve comparisons between elements. We hypothesize that alterations in chart elements can shift GPT-4o’s focus towards visual comparisons rather than numerical ones.
The bar chart on the right in Figure 6 illustrates the varied impacts of 15 chart variants on GPT-4o’s performance. We posit that data labels play a crucial role in GPT-4o’s low-level data analysis capabilities, as removing or reducing them tends to diminish its effectiveness. Furthermore, adding marks to the legend or eliminating the legend’s color significantly affects GPT-4o’s performance on specific tasks by introducing visual clutter or removing essential visual cues, respectively. For instance, changes in color and legend can greatly assist GPT-4o in solving Correlation tasks.
Finding-9: While most chart variants have a minimal impact on GPT-4o’s performance, the absence of data labels significantly affects its accuracy. Additionally, larger labels and the removal of data labels can actually enhance GPT-4o’s performance in anomaly detection and filtering tasks, as it shifts its focus to visual comparisons.
◆ Exp-6: Impact of Image Quality
Experimental Settings.
In this experiment, we evaluate the robustness of GPT-4o in low-level ChartQA tasks by introducing six types of noise, as shown in Figure 8 (b). We use 245 visual variants of 35 charts as testing samples, along with 8,456 textual prompts covering 10 tasks.
Overall Results.
Figure 21 reports the experimental results. Generally, six methods of degrading image quality tend to negatively impact GPT-4o across a broad range of tasks and chart types. Among these, Median Blur stands out as the most detrimental, causing an average performance decline of 16.8%. We consider that median blurring makes numerical labels unreadable, resulting in a significant decrease in the performance of tasks directly related to numerical values. Interestingly, both increasing and decreasing the brightness show positive effects on the majority of tasks, with an average improvement of 0.6% and 1.3% respectively. In addition, Figure 21(b) reveals a unique finding where the distribution and cluster tasks did not experience any negative impact under six types of noise. In fact, the performance of the distribution task improved by an average of 9.17%, while the cluster task improved by an average of 5.28%.
Finding-10: GPT-4o’s accuracy varies under the influence of different types of noise on various charts. Interestingly, there are instances where the accuracy improves, particularly in visually semantic tasks. This suggests that GPT-4o can rely more on visual information when the textual information is compromised.
@@
Table 4: GPT-4o: The Synergistic Effect of Visual and Textual Prompts (Overall Accuracy (%))
@@
Basic Textual Prompts 55.9 34.0 70.1 68.5 80.6 68.9 49.9 82.6 93.9 74.3 69.2
ChartCoT 75.7 72.5 65.0 63.7 80.0 62.5 60.0 89.2 92.5 87.5 76.1
Role-Play 61.8 72.5 60.0 71.2 85.8 71.2 64.2 87.5 96.7 80.0 74.6
Tutorial 75.0 77.5 80.0 81.2 86.7 37.5 75.8 86.7 94.2 90.0 78.4
Chain-of-Charts (ours) 78.9 52.5 75.0 71.3 89.2 86.3 76.7 96.7 97.5 92.5 83.5
@@
Basic Textual Prompts 77.5 62.5 60.0 68.8 89.2 62.5 63.3 95.0 98.3 90.0 79.4
ChartCoT 80.7 65.0 55.0 71.2 85.0 56.2 60.8 91.7 92.5 87.5 78.0
Role-Play 77.9 70.0 55.0 75.0 90.8 55.0 61.7 94.2 99.2 85.0 79.4
Tutorial 80.0 80.0 75.0 85.0 92.5 32.5 75.8 90.8 99.2 92.5 81.6
Chain-of-Charts (ours) 83.2 62.5 65.0 65.0 90.8 77.5 77.5 96.7 99.2 95.0 84.3
◆ Exp-7: Synergistic Effect of Question Types, Textual Prompts, and Visual Prompts
Experimental Settings.
As previously mentioned, using Visual Prompts alone cannot robustly improve the accuracy of GPT-4o. By comparing the bar charts on the right side of Figure 5(a) and (c), we can see that the accuracy of GPT-4o using Visual Prompts actually decreases in Distribution and Order tasks. In contrast, by comparing the bar charts on the right side of Figure 5(a) and (c), we find that under the influence of Chain-of-Charts, GPT-4o has made significant improvements in all 10 tasks. Therefore, we ask: Can the combination of visual and text prompts enhance performance in low-level ChartQA tasks with GPT-4o? We use the same samples as in Exp-4 for this experiment.
Overall Results.
Figure 5(d) and shows GPT-4o’s accuracy following the integration of Chain-of-Charts and Visual Prompt, demonstrating a clear enhancement over the outcomes depicted in Figures 5(a), (b), and (c), which demonstrates the combined strategy’s effectiveness.
Table 4 reports the performance improvements in GPT-4o after integrating various textual prompts with visual prompts. Furthermore, this combination attained the highest accuracy in six tasks.
@@
As depicted in Figure 22, the integration of Chain-of-Charts and visual prompts enables GPT-4o to outperform other settings. However, the improvement over using Chain-of-Charts alone is slight. We discuss the possible reasons behind:
@@
First, after carefully analyzing the experimental results, we discover that GPT-4o exhibits a certain degree of hallucination in chart understanding. For example, even if the calculation process is accurate, GPT-4o may provide answers that do not match any of the multiple-choice options, leading to incorrect results. This indicates that the model’s accuracy is significantly affected by hallucination. Moreover, we also observe that GPT-4o might output numerical information unrelated to the chart even when explicitly recognizing values, further evidencing the hallucination phenomenon in chart reading.
Finding-12: Adding a Visual Prompt improves performance, but its impact is limited when applied to the Chain-of-Charts strategy.
4 Related Work
Low-Level ChartQA Tasks.
Low-level data analysis tasks in chart involve activities such as data retrieval and correlation determination. These tasks were defined by Amar et al. (2005) and later evaluated by Saket et al. (2019) in a crowdsourced experiment. In this paper, we use these ten tasks as a framework to assess the effectiveness of MLLMs in low-level ChartQA.
Evaluating MLLMs in ChartQA Tasks.
Recent studies Cheng et al. (2023); Zhou et al. (2023); Huang et al. (2023); Xu et al. (2023); Han et al. (2023); Xia et al. (2024); Masry et al. (2023) have leveraged MLLMs to perform high-level ChartQA tasks, such as chart captioning. For example, Huang et al. (2023) evaluated the capabilities of representative MLLMs, such as GPT-4V and Gemini Team et al. (2024), on chart captioning tasks, finding challenges in accurately reflecting factual chart information. Diverging from this focus on high-level tasks, our research uniquely targets low-level ChartQA tasks Saket et al. (2019).
ChartQA Datasets.
In the last decade, several ChartQA datasets have been presented, as shown in Table 1. For example, ChartBench Xu et al. (2023) includes 2.1K charts for four types of ChartQA tasks. However, a gap remains - no existing dataset comprehensively evaluates the 10 critical low-level ChartQA tasks Amar et al. (2005). In addition, to conduct customized evaluations, we need access to chart metadata (e.g., underlying data), not just images. Therefore, we curate a large-scale dataset, ChartInsights, consisting of 22,347 quartets - each with a chart, a query, and its answer.
We also include more detailed discussion about the related work in Appendix E.